# Perturbation Theory and the Quantum Rabi-model

### v1

@@ 4.1. Study of $(T_{+}\varphi,T_{-}\varphi)_{0}$
(by the Taylor expansion of Hermite polynomials)
@@
In other words, Lemma 4.2 makes sure that if $x_{0}=4\alpha^{2}\in(0,+\infty)\setminus\mathsf{Z}$ then given any $j\geq 1$ we may find $k_{j}\geq 1$ such that for all integers $N\leq k_{j}$ the Laguerre polynomials $L_{N}$ do not have zeros at distance less than $\delta_{j}.$ Hence, by Lemma 4.1 for all such $N$ we have that $(T_{+}\varphi,T_{-}\varphi)_{0}\not=0$, whence (as we will see through Rellich’s theory) the perturbations of $\lambda_{N}(0)$ (which has multiplicity 2) split into simple eigenvalues $\lambda_{N}^{\pm}(\varepsilon)$. Recall, in fact, that
@@
(convergence of the series in $L^{2}$) with $u^{(N)}_{\pm}(\varepsilon)\in D(A+\varepsilon B)$ for all $|\varepsilon|<c_{j}$ and
@@
By unitary conjugation we therefore obtain also the following result.
@@
of $Q_{R,\varepsilon}$ are all simple.
We will call the $\lambda^{(N)}_{\pm}(\varepsilon)=\mu^{(N)}_{\pm}(\varepsilon)+\alpha^{2}/2$ the shifted eigenvalues of $Q_{R}$.
@@ 5. The Braak conjecture
In [1], Braak conjectured that, within an interval $I_{N}:=[N,N+1),$ $N\in\mathbb{Z}_{+}$, there may be lying either $0,$ or $1$ or at most $2$ shifted eigenvalues $\lambda+g^{2}=\lambda+\alpha^{2}/2$, $\lambda\in\mathrm{Spec}(Q_{\mathrm{Rabi}})$, of the Rabi system. Moreover, if in $I_{N}$ there are two roots of the function $G_{+}$ (resp. $G_{-}$; see [1]) then the intervals $I_{N-1}$ or $I_{N+1}$ may contain zero or at most $1$ root of $G_{+}$ (analogously for $G_{-}$). Otherwise said (see [22]), Braak’s $G$-conjecture states that for a given parity (given by $N$ being even or odd; the system preserves parity so that the even part and the odd part of $L^{2}$ are invariant and orthogonal) the interval $I_{N}$ contains at most two shifted eigenvalues, two intervals containing no shifted eigenvalues are not adjacent, and two intervals containing two shifted eigenvalues are also not adjacent. The latter conjecture has been proved by Rudnik in [22] ”for almost all $N$s”.
In our case, that is the system $Q_{R}$ (a particular Rabi system, due to the condition on the interaction coefficient $\alpha$ and the zeroth-order term in the symbol $Q_{R}(X)$), we have that by Theorem 4.3 for any given $j\geq 1$ there is a $k_{j}\in\mathbb{Z}_{+}$ and $c_{j}>0$ such that for the first $2(k_{j}+1)$ shifted eigenvalues of $Q_{R}$ , the $\lambda^{(N)}_{\pm}(\varepsilon)$ with $N\leq k_{j},$ we have, for $|\varepsilon|<c_{j}$,
(5.1) $\lambda_{N,\pm}(\varepsilon)=N+\frac{1}{2}+\mu_{\pm}\varepsilon+o(\varepsilon)\in I_{N},\quad N=0,1,\ldots,k_{j},$
@@
We therefore have the following corollary.
Corollary 5.1.
Let $\alpha\in\mathbb{R}$ be such that $4\alpha^{2}\in(0,+\infty)\setminus\mathsf{Z}.$ For the system $Q_{R}$, Braak’s conjecture holds for every spectral initial segment of eigenvalues: For any given $j\geq 1$ there exists $k_{j}\in\mathbb{Z}_{+}$ ($k_{j}\nearrow+\infty$) and $c_{j}>0$ sufficiently small such that the conjecture holds for the eigenvalues belonging to
$\mathrm{Spec}(Q_{R})\cap\Bigl(I_{N}-\alpha^{2}/2\Bigr),\quad N=0,\ldots,k_{j},$
for all $|\varepsilon|<c_{j}.$
6. The case in which $4\alpha^{2}\in\mathsf{Z}$
In this section we tackle the case in which $\alpha$ is such that $4\alpha^{2}\in\mathsf{Z}$, that is, we have a family (possibly infinite) of Laguerre polynomials $L_{k}$. So, we let
$\mathsf{K}:=\{k\in\mathbb{Z}_{+};\,\,L_{k}(4\alpha^{2})=0\}.$
@@ 6. The case in which $4\alpha^{2}\in\mathsf{Z}$
With $k\in\mathsf{K}$ we look for
@@
For $N\in\mathsf{K}$, let $\mu_{+}=(\gamma_{1}+\gamma_{2})/2$ and let $Q^{(2)}_{N}\colon\mathbb{C}^{2}\times\mathbb{C}^{2}\longrightarrow\mathbb{C}$ be the sesquilinear form
@@
Let $N\in\mathsf{K}.$ Let $Q^{(2)}_{N}$ be the quadratic form (6.3), that we suppose to possess two distinct eigenvalues $\mu_{\pm,N}^{(2)}$ (necessarily real). Let $w_{\pm}\in\mathbb{C}^{2}$ be corresponding orthonormal vectors belonging to $\mu^{(2)}_{\pm,N}$, respectively. Then, putting in (6.2)
@@
is a family of quasi-modes (see [10]) such that
@@
By [10], we then have that there must be an eigenvalue of $A+\varepsilon B$ within each of the $\pm$-intervals
@@
But since the eigenvalues are analytic in $\varepsilon$ and the multiplicity of the eigenvalue, for $\varepsilon$ small, is at most $2$ we therefore have that also in this case near $N+1/2$ there are exactly $2$ distinct eigenvalues. By unitary conjugation we therefore have the following theorem.
@@
Suppose $4\alpha^{2}\in\mathsf{Z}$. For each $N\in\mathsf{K}$ for which the quadratic form $Q^{(2)}_{N}$ over $\mathbb{C}^{2}$ has distinct eigenvalues, we have that $Q_{R}$, for $\varepsilon\not=0$ sufficiently small, has exactly $2$ eigenvalues near $N+1/2$.
Consequently, Braak’s conjecture holds also for these $N$s.
@@
We wish in this subsection to show how it is possible to explicitly write down the quadratic form $Q_{N}^{(2)}.$ It shows that a combination of Rellich’s formulae and a quasimode construction gives the possibility to study highly degenerate eigenvlues. Since as functions of $\varepsilon$ the eigenvalues are analytic, this gives the opportunity to understand the spectrum (of course, in the perturbative regime) of the QRM.

### v2

@@ 4.1. Study of $(T_{+}\varphi,T_{-}\varphi)_{0}$
(by the Taylor expansion of Hermite polynomials (3.4))
@@
In other words, Lemma 4.2 makes sure that if $x_{0}=4\alpha^{2}\in(0,+\infty)\setminus\mathsf{Z}$ then given any $j\geq 1$ we may find $k_{j}\geq 1$ such that for all integers $N\leq k_{j}$ the Laguerre polynomials $L_{N}$ do not have zeros at distance less than $\delta_{j}.$ Hence, by Lemma 4.1 for all such $N$ we have that $(T_{+}\varphi,T_{-}\varphi)_{0}\not=0$, whence (as we will see through Rellich’s theory) the perturbations of $\lambda_{N}(0)$ (which has multiplicity 2) split into simple eigenvalues $\lambda^{(N)}_{\pm}(\varepsilon)$. Recall, in fact, that
@@
with $u^{(N)}_{\pm}(\varepsilon)\in D(A+\varepsilon B)$ for all $|\varepsilon|<c_{j}$ and
@@
By unitary conjugation we therefore obtain also the following result for the eigenvalues of our Rabi quantum model $Q_{R}$.
@@
of $Q_{R,\varepsilon}$ are all simple when $\varepsilon\not=0$.
We will call the $\lambda^{(N)}_{\pm}(\varepsilon)=\mu^{(N)}_{\pm}(\varepsilon)+\alpha^{2}/2$ the shifted eigenvalues of $Q_{R,\varepsilon}$.
Recall, at this point, the parity associated with Rabi’s system $Q_{\mathrm{Rabi},\varepsilon}$ (but also with our system $Q_{R,\varepsilon}$), given by the parity operator
$\Pi=(-1)^{\boldsymbol{a}^{\dagger}\boldsymbol{a}}\sigma_{z}.$
Since $\Pi$ commutes with $Q_{R}$ we have an orthogonal splitting
$L^{2}(\mathbb{R};\mathbb{C}^{2})=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-},$
where $\boldsymbol{H}_{\pm}$ are the eigenspaces of the parity operator belonging to the eigenvalues $\pm 1$, respectively. Recall that the structure of $\boldsymbol{H}_{\pm}$ is the following:
$\boldsymbol{H}_{+}=\left\{\left[\begin{array}[]{c}f_{1}\\ f_{2}\end{array}\right]\in L^{2}(\mathbb{R};\mathbb{C}^{2});\,\,\text{\rm$f_{1}$ is even and $f_{2}$ is odd}\right\},$
$\boldsymbol{H}_{-}=\left\{\left[\begin{array}[]{c}f_{1}\\ f_{2}\end{array}\right]\in L^{2}(\mathbb{R};\mathbb{C}^{2});\,\,\text{\rm$f_{1}$ is odd and $f_{2}$ is even}\right\}.$
Now, Theorem 4.3 gives also a precise control on the parity of the eigenvalues of $Q_{R}$ by virtue of the control on the analytic family of eigenvectors (which are the conjugation of the eigenfunctions $u_{\pm}^{(N)}(\varepsilon)$ via the metaplectic operator $U$ in Lemma 3.6 ). We start from the case $\varepsilon=0$.
Let $v_{\pm}^{(N)}(\varepsilon):=U_{0}T^{*}u_{\pm}^{(N)}(\varepsilon)$, $|\varepsilon|<\varepsilon_{0}$, be the eigenfunctions associated with the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ of $Q_{R,\varepsilon}$, respectively, which are simple if $\varepsilon\neq 0$. Also, put for short
$s_{N}:=\mathrm{sgn}(T_{+}\varphi,T_{-}\varphi)_{0}\quad\text{\rm and}\quad\sigma_{N}=(-1)^{N}s_{N}.$
Lemma 4.5.
In the notation and with the hypotheses of Theorem 4.3, the eigenvalue $\lambda_{+}^{(N)}(0)=\lambda_{-}^{(N)}(0)=N+\frac{1}{2}$ of $Q_{R,0}$ belongs both to $\mathrm{Spec}\,(Q_{R,0}|_{\boldsymbol{H}_{+}})$ and to $\mathrm{Spec}\,(Q_{R,0}|_{\boldsymbol{H}_{-}})$. In fact $v_{\pm}^{(N)}(0)\in\boldsymbol{H}_{\pm\sigma_{N}}$.
Proof.
We have that $N+\frac{1}{2}$ has multiplicity $2$ with $v_{\pm}^{(N)}(0)=U_{0}T^{*}u_{\pm}^{(N)}(0)$ generating the eigenspace of $Q_{R,0}$ associated with $N+\frac{1}{2}$ by Theorem 4.3.
If, say, $v_{+}^{(N)}(0)$ had nontrivial projections onto $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{-}$, then also $v_{+}^{(N)}(\varepsilon)$ would have nontrivial projections onto $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{+}$, for small $\varepsilon,$ whence by orthogonality the multiplicity of $\lambda_{+}^{(N)}(\varepsilon)$ would be 2 for $\varepsilon\not=0$ and small, which is impossible. Therefore, $\pm$-respectively, $v_{\pm}^{(N)}(0)$ belongs to either $\boldsymbol{H}_{+}$ or $\boldsymbol{H}_{-}$.
Next, a direct computation gives
$v_{\pm}^{(N)}(0)=\frac{1}{\|\varphi\|_{0}\sqrt{2}}\begin{bmatrix}s_{N}T_{-}\varphi\pm T_{+}\varphi\\ s_{N}T_{-}\varphi\mp T_{+}\varphi\end{bmatrix}$
and we study when, $\pm$-respectively, $v_{\pm}^{(N)}(0)$ belongs to $\boldsymbol{H}_{+}$, the case $\boldsymbol{H}_{-}$ being analogous. It is possible if and only if
$s_{N}T_{-}\varphi\pm T_{+}\varphi\quad\text{\rm is even and}\,\,s_{N}T_{-}\varphi\mp T_{+}\varphi\quad\text{\rm is odd.}$
This implies
$s_{N}\varphi(x+\alpha)=\pm\varphi(-x-\alpha),\forall x\in\mathbb{R},$
which is equivalent to
$s_{N}\varphi(x+\alpha)=\pm(-1)^{N}\varphi(x+\alpha),\forall x\in\mathbb{R},$
since an Hermite polynomial is even or odd according to whether its degree is even or odd, respectively. This proves the claim. ∎
Now we generalize the result to $\varepsilon\neq 0$.
Lemma 4.6.
In the notation and with the hypotheses of Theorem 4.3, the eigenvalue $\lambda_{\pm}^{(N)}(\varepsilon)$ belongs to $\mathrm{Spec}\,$($Q_{R,0}|_{\boldsymbol{H}_{\pm\sigma_{N}}}$)  (that is, it has parity $\pm\sigma_{N}$) for all $|\varepsilon|<\varepsilon_{0}$.
Proof.
[…]
@@ 5. The Braak conjecture
In [1], Braak conjectured that, within an interval $I_{N}:=[N,N+1),$ $N\in\mathbb{Z}_{+}$, there may be lying either $0,$ or $1$ or at most $2$ shifted eigenvalues $\lambda+g^{2}=\lambda+\alpha^{2}/2$, $\lambda\in\mathrm{Spec}(Q_{\mathrm{Rabi}})$, of the Rabi system. Moreover, if in $I_{N}$ there are two roots of the function $G_{+}$ (resp. $G_{-}$; see [1]) then the intervals $I_{N-1}$ or $I_{N+1}$ may contain zero or at most $1$ root of $G_{+}$ (analogously for $G_{-}$). Otherwise said (see [23]), Braak’s $G$-conjecture states the following: For each parity $N$ determined by the parity operator $\Pi$, the interval $I_{N}$ contains at most two shifted eigenvalues, two intervals containing no shifted eigenvalues are not adjacent, and two intervals containing two shifted eigenvalues are also not adjacent.
The latter conjecture has been proved by Rudnik in [23] ”for almost all $N$s”.
It is also convenient to restate Braak’s conjecture as in Lanuza [10]: Let $\varepsilon\not=0$ and let $B_{N}^{\pm}=\#\>\mathrm{Spec}\,(Q_{\mathrm{Rabi},\varepsilon}|_{\boldsymbol{H}_{\pm}})\cap\bigl[N+1/2,N+3/2\bigr)$. For all $N\geq 0$ and each choice of $\pm$, we have that $B_{N}^{\pm}\in\{0,1,2\}$ and $B_{N}^{\pm}+B_{N+1}^{\pm}\in\{1,2,3\}$.
In addition, as we shall see, we may state: (for $\varepsilon\not=0$ the interval $\bigl[-1/2,1/2\bigr)$ contains exactly one eigenvalue of parity $\sigma_{0}$.
Recall that for the system $Q_{R,\varepsilon}$, we have by Theorem 4.3 that for any given $j\geq 1$ there is a $k_{j}\in\mathbb{Z}_{+}$ and $c_{j}>0$ such that for the first $2(k_{j}+1)$ shifted eigenvalues of $Q_{R,\varepsilon}$ , the $\lambda^{(N)}_{\pm}(\varepsilon)$ with $N\leq k_{j},$ we have, for $|\varepsilon|<c_{j}$,
(5.1) $\lambda_{N,\pm}(\varepsilon)=N+\frac{1}{2}+\mu_{\pm}\varepsilon+o(\varepsilon),\quad N=0,1,\ldots,k_{j},$
@@
In view of Theorem 4.3, for $Q_{\mathrm{Rabi},\varepsilon}$ (and $\varepsilon\not=0$) we have the following corollary, which proves Braak’s Conjecture for finite segments of the spectrum. Because of the shift by $1/2$ in the spectrum, due to our choice of definition of $Q_{R,\varepsilon}$, the intervals in which we have to test Braak’s conjecture for the $\lambda_{\pm}^{(N)}(\varepsilon)$s are given by
$I_{N+1/2}:=\bigl[N+1/2,N+3/2\bigr),\,\,\,\,N\in\mathbb{Z}_{+}.$
Corollary 5.1 (Braak’s Conjecture).
If $4\alpha^{2}\not\in\mathsf{Z}$ and $\gamma_{1}=-\gamma_{2}$, the Braak conjecture holds for every spectral initial segment of eigenvalues: For any given $j\geq 1$ there exist $k_{j}\in\mathbb{Z}_{+}$ ($k_{j}\nearrow+\infty$) and $c_{j}>0$ sufficiently small such that the conjecture holds for the interval $I_{N+1/2}$ for all $|\varepsilon|<c_{j}$ and all $N<k_{j}-2$. More precisely, when $\varepsilon\not=0$ for each interval $\bigl[N+1/2,N+3/2\bigr)$ ($N\geq 0$) there are exactly two simple eigenvalues of $Q_{\mathrm{Rabi},\varepsilon}$ with parity $\sigma_{N}$ and $\sigma_{N+1}$, respectively. When $N=0$ (and $\varepsilon\not=0$) in $\bigl[-1/2,1/2\bigr)$ there is exactly one simple eigenvalue with parity $\sigma_{0}$.
Proof.
When $Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}+\frac{1}{2}I_{2}$, one has $\gamma_{1}=-\gamma_{2}$ and get that $\beta_{1}=0$, $\beta_{2}=\gamma_{1}$ and, hence, $\pm\mu_{\pm}>0$ (same choice of $\pm$). Hence, for all $0\leq N<k_{j}-2$ and for $|\varepsilon|\not=0$ sufficiently small, $\lambda_{+}^{(N)}(\varepsilon),\,\lambda_{-}^{(N+1)}(\varepsilon)\in\bigl[N+1/2,N+3/2\bigr)$ and $\lambda_{+}^{(N+1)}(\varepsilon),\,\lambda_{-}^{(N+2)}(\varepsilon)\in\bigl[N+3/2,N+5/2\bigr)$ and these eigenvalues are simple by Theorem 4.3. Hence, for such $N$s, $B_{N}^{\pm}\leq 2$. Moreover, by Lemma 4.6, $\lambda_{+}^{(N)}(\varepsilon)$, $\lambda_{-}^{(N+1)}(\varepsilon)$, $\lambda_{+}^{(N+1)}(\varepsilon)$, and $\lambda_{-}^{(N+2)}(\varepsilon)$ have parity $\sigma_{N}$, $\sigma_{N+1}$, $\sigma_{N+1}$, and $\sigma_{N+2}$, respectively. Hence, as $\sigma_{N+1}$ and $\sigma_{N+1}$ have different sign, and since there are four eigenvalues in the union $\bigl[N+1/2,N+3/2\bigr)\cup\bigl[N+3/2,N+5/2\bigr)$ and they are all simple, we proved $1\leq B_{N+1}^{\pm}+B_{N+2}^{\pm}\leq 3$.
Hence, Braak’s conjecture holds. ∎
Therefore, Braak’s conjecture may fail just when we are in a degenerate case, that is when $4\alpha^{2}$ is a zero of at least a Laguerre polynomial. This case will be the topic of next section.
6. The case $4\alpha^{2}\in\mathsf{Z}$
In this section we tackle the case in which $\alpha$ is such that $4\alpha^{2}\in\mathsf{Z}$, that is, we have a family (possibly infinite) of Laguerre polynomials $L_{k}$ that vanish at $4\alpha^{2}$. So, we let
$\mathsf{K}_{\alpha}:=\{k\in\mathbb{Z}_{+};\,\,L_{k}(4\alpha^{2})=0\}.$
Hence $(T_{+}\varphi_{k},T_{-}\varphi_{k})_{0}=0$ for all $k\in\mathsf{K}_{\alpha}$ and Rellich’s approach fails in this case.
Remark 6.1.
Because of the interlacing property of the zeros of Laguerre’s polynomial we have
$k\in\mathsf{K}_{\alpha}\Longrightarrow k-1,k+1\not\in\mathsf{K}_{\alpha}.$
@@ 6. The case $4\alpha^{2}\in\mathsf{Z}$
With $k\in\mathsf{K}_{\alpha}$ we look for
@@
For $N\in\mathsf{K}_{\alpha}$, let $\mu_{+}=(\gamma_{1}+\gamma_{2})/2$ and let $Q^{(2)}_{N}\colon\mathbb{C}^{2}\times\mathbb{C}^{2}\longrightarrow\mathbb{C}$ be the sesquilinear form
@@
Let $N\in\mathsf{K}_{\alpha}.$ Let $Q^{(2)}_{N}$ be the bilinear form (6.3), that we suppose to possess two distinct eigenvalues $\mu_{\pm,N}^{(2)}$ (necessarily real). Let $w_{\pm}\in\mathbb{C}^{2}$ be corresponding orthonormal vectors belonging to $\mu^{(2)}_{\pm,N}$, respectively. Put in (6.2)
@@
is a family of quasi-modes (in the sense of Lazutkin, see [11]) such that
@@
By [11], there must then be an eigenvalue of $A+\varepsilon B$ within each of the $\pm$-intervals
@@
But since the eigenvalues are analytic in $\varepsilon$ and the multiplicity of the eigenvalue, for $\varepsilon$ small, is at most $2$ (since the projector $\varepsilon\mapsto(2\pi i)^{-1}\int_{\Gamma}(Q_{R,\varepsilon}-z)^{-1}dz$, with $\Gamma=\{|z|=\delta\}$ contained in the resolvent set of $Q_{R,0}$, which is the projector onto the eigenspace of $N+1/2$ for $\varepsilon=0$, is holomorphic for small $\varepsilon$) we therefore have that also in this case near $N+1/2$ there are exactly $2$ distinct eigenvalues. By unitary conjugation we therefore have the following theorem.
@@
Suppose $4\alpha^{2}\in\mathsf{Z}$. For each $N\in\mathsf{K}_{\alpha}$ for which the quadratic form $Q^{(2)}_{N}$ over $\mathbb{C}^{2}$ has distinct eigenvalues, we have that $Q_{R,\varepsilon}$, for $\varepsilon\not=0$ sufficiently small, has exactly $2$ eigenvalues near $N+1/2$.
When considering QRM’s spectrum near $N+1/2$ for $N\in\mathsf{K}_{\alpha}$, we have that $\mu_{+}=\mu_{-}=0,$ so that the unperturbed degenerate eigenvalue $\lambda^{(N)}_{\pm}(0)=N+1/2$ is subjected to an $\varepsilon^{2}$ perturbation. Hence, depending on whether $Q_{N}^{(2)}$ is nondegenerate and is >0>0, resp. <0<0 definite, using Remark 6.1 and the analysis of the last section, in an interval $\bigl[N+1/2,N+3/2)$ we may find, to fix ideas in case $Q_{N}^{(2)}>0$ say, the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ and $\lambda_{-}^{(N+1)}(\varepsilon)$ and in this case have $0\leq B_{N}^{\pm}\leq 3$ while $0\leq B_{N-1}^{\pm}\leq 1$. However, when $Q_{N}^{(2)}$ is nondegenerate with signature $(1,1)$, one may again have a configuration of the spectrum consistent with Braak’s conjecture.
The point is to control the parity of the eigenfunctions near the eigenvalue $N+1/2$ of the unperturbed operator.
Lemma 6.5.
For all $N\geq 0$ and for any choice of $\pm$, $\Pi_{\boldsymbol{H}_{\pm}}\mathrm{Ker}(Q_{R,0}-(N+1/2))\neq\{0\}$ where $\Pi_{\boldsymbol{H}_{\pm}}$ is the projection onto $\boldsymbol{H}_{\pm}$.
Proof.
In the first place, we know that
(6.4) $\mathrm{Ker}(Q_{R,0}-(N+1/2))=\mathrm{Span}\,\{U_{0}T^{*}(\phi_{N}\otimes e_{\pm})\}=\mathrm{Span}\,\Big\{(T_{\mp}\phi_{N})\otimes\begin{bmatrix}1\\ \pm 1\end{bmatrix}\Big\}$,
where $e_{+}=\left[\begin{array}[]{c}1\\ 0\end{array}\right]$ and $e_{-}=\left[\begin{array}[]{c}0\\ 1\end{array}\right]$, respectively. Next, since $\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}=L^{2}$, it is sufficient to show that $\mathrm{Ker}(Q_{R,0}-(N+1/2))\not\subset\boldsymbol{H}_{\pm}$ to prove the claim. In fact, by (6.4) having $\mathrm{Ker}(Q_{R,0}-(N+1/2))\subset\boldsymbol{H}_{+}$ is equivalent to $T_{\mp}\phi_{N}$ is even and $\pm T_{\mp}\phi_{N}$ is odd. However, this is impossible since the only function being both even and odd is 0. The same argument applies in case of inclusion in $\boldsymbol{H}_{-}$. This concludes the proof. ∎
The foregoing lemma allows us to obtain a result on the parity of the two distinct eigenvalues in Theorem 6.4 when $\varepsilon\neq 0$, as the next result shows.
Lemma 6.6.
In the notation of Theorem 6.4, the two distinct eigenvalues near $N+1/2$ have different parity for $\varepsilon\neq 0$.
Proof.
Let $\varepsilon\mapsto v_{\pm}^{(N)}(\varepsilon)$ be the analytic family of eigenfunctions of $Q_{R,\varepsilon}$ associated with the eigenvalues $\lambda_{\pm}^{(N)}(\varepsilon)$ of $Q_{R,\varepsilon}$.
As a first step, we show that either $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{+}$ or $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{-}$ (we can proceed similarly for $v_{-}^{(N)}(\varepsilon)$) for all $\varepsilon\neq 0$. In fact, $\Pi_{\boldsymbol{H}_{\pm}}v_{+}^{(N)}(\varepsilon)$ also satisfy the eigenvalue equation with respect to $\lambda_{+}^{(N)}(\varepsilon),$ since $Q_{R,\varepsilon}=Q_{\mathrm{Rabi},\varepsilon}$ (recall that $\gamma_{1}=-\gamma_{2}$) commutes with the parity operator. If both projections were nonzero, they would be eigenfunctions as well, and because of their orthogonality they would be linearly independent, yielding that $\lambda_{+}^{(N)}(\varepsilon)$ could not be simple for $\varepsilon\not=0$.
Next, as a second step, we show that if $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{\sigma}$ ($\sigma=1$ or 1-1) for some $\varepsilon$, then $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{\sigma}$ for all $\varepsilon$ (we can proceed similarly for $v_{-}^{(N)}$). Assume, to fix ideas, $v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}$. If there were $\varepsilon^{\prime}\not=0$ and $\delta>0$ such that $v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}$ for $\varepsilon\in(\varepsilon^{\prime}-\delta,\varepsilon^{\prime})$ and $v_{+}^{(N)}(\varepsilon)\in\boldsymbol{H}_{-}$ for $\varepsilon\in(\varepsilon^{\prime},\varepsilon^{\prime}+\delta)$, then by continuity of the family $\varepsilon\mapsto v_{+}^{(N)}(\varepsilon)$ we would have $0\neq v_{+}^{(N)}(\varepsilon^{\prime})\in\boldsymbol{H}_{+}\cap\boldsymbol{H}_{-}=\{0\}$ (by the orthogonality of $\boldsymbol{H}_{+}$ and $\boldsymbol{H}_{-}$).
As a final step, we show that the two eigenvalues $v_{\pm}^{(N)}(\varepsilon)$ cannot be in the same space $\boldsymbol{H}_{+}$ or $\boldsymbol{H}_{-}$ for any fixed $\varepsilon$. In fact, if by contradiction we had that both eigenfunctions $v_{\pm}^{(N)}(\varepsilon)\in\boldsymbol{H}_{+}$ for some $\varepsilon$ (we can proceed similarly for $\boldsymbol{H}_{-}$), then, by the second step, we would have $v_{\pm}^{(N)}(0)\in\boldsymbol{H}_{+}$ and, hence, $\Pi_{\boldsymbol{H}_{-}}\mathrm{Ker}(Q_{R,0}-(N+1/2))=\Pi_{\boldsymbol{H}_{-}}\mathrm{Span}\,\{v_{\pm}^{(N)}(0)\}=\{0\}$ against Lemma 6.5. This concludes the proof. ∎
We may therefore state the following corollary of our approach.
Corollary 6.7 (Braak’s conjecture, degenerate case).
Consider our system $Q_{R,\varepsilon}$ in the case of $Q_{\mathrm{Rabi},\varepsilon}$, i.e. when $\mu_{\pm}=0$. Suppose the bilinear form $Q_{N}^{(2)}$ is nondegenerate with signature $(1,1).$ Then Braak’s conjecture keeps holding true.
It will be interesting also to consider the other cases depending on the positivity of $Q_{N}^{(2)}$, and also cases in which the eigenvalues are either constant (i.e. $N+1/2$ for some $N\in\mathsf{K}_{\alpha}$) or have derivatives with respect to $\varepsilon$ at 0 all zero up to some $k_{0}\in\mathbb{N}$ (hence $d^{k_{0}+1}\lambda_{\pm}^{(N)}/d\varepsilon^{k_{0}+1}(0)\not=0$). In that case the behavior will be decided (using, for instance, a quasi-mode construction) by an analogous bilinear form similar to $Q_{N}^{(2)}$ in the case $k_{0}=1$.
@@
In this subsection we wish to show how it is possible to explicitly write down the quadratic form $Q_{N}^{(2)}.$ It shows that a combination of Rellich’s formulae and a quasimode construction gives the possibility to study highly degenerate eigenvalues. Since as functions of $\varepsilon$ the eigenvalues are analytic, this gives the opportunity to understand the spectrum (of course, in the perturbative regime) of the QRM.